# Intensity doubling for Brownian loop-soups in high dimensions

### v2

@@ 4.2. Statement of the key proposition and consequences
So, together with Corollary 8, this proves a number of the statements in Theorem 1 – the remaining to be proved ones being essnetially the ones about the presence/absence of mesoscopic loops that we will derive in Section 4.4.
@@ 5. Concluding remarks
    For each given Brownian loop $B$ in $\Lambda_{N}$, one can sample an independent loop-soup in $\Lambda_{N}$ and then look at the loop-soup cluster $c(B)$ containing $B$ of the overlay of $B$ with this loop-soup. If $B$ is defined according to the Brownian loop-measure, then this procedure defines a measure on clusters $c(B)$. Similarly, for any fixed $\varepsilon$, if $B$ is defined according to the Brownian loop-measure restricted to ${\mathcal{B}}(\varepsilon,N)$, this defines a measure $\rho_{\varepsilon}$ on the set of clusters in $\Lambda_{N}$. The fact that (with probability that goes to $1$ as $N\to\infty$) no macroscopic loop-cluster contains more than one loop in ${\mathcal{B}}(\varepsilon,N)$ and dimension considerations readily imply that when $d\geq 9$ (up to an event of probability that vanishes as $N\to\infty$), the collection of clusters that do contain loops in ${\mathcal{B}}(\varepsilon,N)$ is distributed like a Poisson point process with intensity $\rho_{\varepsilon}$ (and that the clusters in this Poisson point process will be disjoint). Combined with our theorem, this actually indicates that the clusters in ${\mathcal{C}}(\varepsilon,N)$ are (asymptotically) distributed like a Poisson point process with intensity $2\rho_{\varepsilon}$, and that all these clusters on cable-graphs will be disjoint.
@@
    Finally, we can note that in the scaling limit (see [10]), rescaled large Brownian loops on the cable-graph $\Lambda_{N}$ converge to continuum Brownian loops in $1,1]^{d}$, so that the theorem implies that the rescaled collection $(\Gamma(C_{1})/N,\ldots,\Gamma(C_{K_{\varepsilon}})/N)$ of cycles of clusters in ${\mathcal{C}}(\varepsilon,N)$ will converge to the corresponding Poissonian collection of Brownian loops (i.e., that wind around some $\Delta$ while staying at distance at least $\varepsilon$ from it), when the intensity of the loop-soup is twice that of the usual one.

### v3

@@ 4.2. Statement of the key proposition and consequences
So, together with Corollary 8, this proves a number of the statements in Theorem 1 – the remaining to be proved ones being essentially the ones about the presence/absence of mesoscopic loops that we will derive in Section 4.4.
@@ 5. Concluding remarks
    For each given Brownian loop $B$ in $\Lambda_{N}$, one can sample an independent loop-soup in $\Lambda_{N}$ and then look at the loop-soup cluster $c(B)$ containing $B$ of the overlay of $B$ with this loop-soup. If $B$ is defined according to the Brownian loop-measure, then this procedure defines a measure on clusters $c(B)$. Similarly, for any fixed $\varepsilon$, if $B$ is defined according to the Brownian loop-measure restricted to ${\mathcal{B}}(\varepsilon,N)$, this defines a measure $\rho_{\varepsilon}$ on the set of clusters in $\Lambda_{N}$. The fact that (with probability that goes to $1$ as $N\to\infty$) no macroscopic loop-cluster contains more than one loop in ${\mathcal{B}}(\varepsilon,N)$ and dimension considerations readily imply that when $d\geq 9$ (up to an event of probability that vanishes as $N\to\infty$), the collection of clusters that do contain loops in ${\mathcal{B}}(\varepsilon,N)$ is distributed like a Poisson point process with intensity $\rho_{\varepsilon}$ (and that the clusters in this Poisson point process will be disjoint). Combined with our theorem, this actually indicates that the clusters in ${\mathcal{C}}(\varepsilon,N)$ are (asymptotically) distributed like a Poisson point process with intensity $2\rho_{\varepsilon}$, and that all these clusters on cable-graphs will be disjoint [mind that we are here discussing about the point process of entire clusters, not just the loops, which explains the $d\geq 9$ constraint].
@@
    We can note that in the scaling limit (see [20]), rescaled large Brownian loops on the cable-graph $\Lambda_{N}$ converge to continuum Brownian loops in $1,1]^{d}$, so that the theorem implies that the rescaled collection $(\Gamma(C_{1})/N,\ldots,\Gamma(C_{K_{\varepsilon}})/N)$ of cycles of clusters in ${\mathcal{C}}(\varepsilon,N)$ will converge to the corresponding Poissonian collection of Brownian loops (i.e., that wind around some $\Delta$ while staying at distance at least $\varepsilon$ from it), when the intensity of the loop-soup is twice that of the usual one.
  9. (9)
    As mentioned in Section 2, instead of working with the loop-soup in $\Lambda_{N}$, one could work with the loop-soup in ${\mathbb{Z}}^{d}$ and consider the collection of all large cycles (i.e., of diameter at least $\varepsilon_{0}N$) that are subsets of $\Lambda_{N}$ and obtain similar results. Let us briefly indicate one way to proceed: The probability that there exists a Brownian loop in the loop-soup that intersects $\Lambda_{N}$ and goes out of $\Lambda_{mN}$ goes to 0 as $m\to\infty$ uniformly with respect to $N$. So, for any very small $\upsilon$, we can first choose $m=m(\upsilon)$ so that this probability is smaller than $\upsilon/2$ for all large $N$, and then note that for $N$ large enough the quantity $u_{mN}$ in Theorem 1 for $\varepsilon=\varepsilon_{0}/m$ is smaller than $\upsilon/2$. This readily shows that large cycles (of diameter at least $\varepsilon N$) within $\Lambda_{N}$ within the whole-space loop-soup clusters correspond either to a chain of small loops (of diameter smaller than $N^{\beta}$) or to the presence of a macroscopic Brownian loop with respective conditional probabilities that are $2\upsilon/2=\upsilon$ close to $1/2$.